OpenStudy (anonymous):

I've been stuck on this for a while now 5-sqrt 20x+4 is greater than or equal to -3

OpenStudy (anonymous):

Please put the brackets in your square root, is it Squareroot (20X+4) or squareroot (20X) +4 ??

OpenStudy (anonymous):

@Karolyn101 squareroot (20x+4)

OpenStudy (anonymous):

5 - sqrt (20x+4) >= -3 -sqrt (20x +4) >= -8 (subtract 5 from both sides) sqrt (20X+4) =< 8 (times -1 both sides, remember to flip the inequality sign) 20X+4 =< 64 (Square both sides) can you solve it now?